# A Gaussian integral formula for the Hermite polynomials: Combinatorics, Asymptotics and Applications

### v1

@@ 6.1 Fourier formula for the Hermite polynomials
The following result corresponds to equation (6.5) in [Sob63].
@@ 6.4 Edgeworth expansion in one dimension
Finally, by the technical assumption that $\left|\phi\right|^{v}$ is integrable, we can use the characteristic/Fourier inversion formula for the density $\rho_{S}(x)=\frac{1}{2\pi}\intop_{-\infty}^{\infty}e^{-\imath ux}\phi_{S_{n}}(u)\mathrm{d}u$ to recover the density from this (see [Fel71] Chapter XV.3 for a detailed proof of this inversion formula). By virtue of the formula for the Hermite polynomials, Proposition 6.4, we can exactly integrate the terms of the approximation for $\phi_{S}$ in (89), as follows:
@@
To complete the proof, it remains only to argue that the $O(\frac{1}{\sqrt{n}^{k-1}})$ error term in the approximation for $\phi_{S}$ in (89) (which holds uniformly for $u$ in a compact set), translates to a $O(\frac{1}{\sqrt{n}^{k-1}})$ error in the density $\rho_{S}$ after integrating with the inversion formula. This error analysis is argued by using the techincal assumption that $\intop_{-\infty}^{\infty}|\phi_{S_{n}}(u)|^{v}\mathrm{d}u<\infty$ which allows one to approximate the inversion formula integral to arbitarty accuracey using $u$ in a compact set; we refer to [Fel71] Chapter XVI.2 for the detailed verification of this error analysis. ∎
@@
Instead of obtaining just the leading order correction, one can obtain an entire series expansion in powers of $n$ known as the Edgeworth expansion. This is achieved by simplifying using more terms in the approximation to the characteristic function (i.e. more cumulants) and also more terms in in the exponential approximation. For example, let us take $k=3$ so that the leading order correction is $\frac{\kappa_{3}}{3!\sqrt{n}}H\!\!\operatorname{e}_{3}(\frac{x}{\sigma})$. By expanding up to order $u^{4}$ and using the higher order correction $\exp(x)=1+x+\frac{1}{2}x^{2}+O(x^{3})$ , one will find that the next order correction is $\frac{1}{n}\left(\frac{\kappa_{4}}{4!\sigma^{4}}H\!\!\operatorname{e}_{4}(\frac{x}{\sigma})+\frac{1}{2}\big{(}\frac{\kappa_{3}}{3!\sigma^{3}}\big{)}^{2}H\!\!\operatorname{e}_{6}(\frac{x}{\sigma})\right)$ where the $H\!\!\operatorname{e}_{6}$ comes from the square term in the exponential which gives an $\frac{1}{2}(\frac{1}{3!}(\imath z)^{3})^{2}$ in the series expansion for $\phi_{S}$. We refer to [Fel71] Chapter XVI for a detailed exposition.
@@
The quantity $S_{n}=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}W_{i}\sqrt{2}\operatorname{ReLU}\left(G_{i}\right)$ where $G_{i},W_{i}$ are iid Gaussians and $\text{ReLU}(x)=\max\{x,0\}$ arises in the theory of neural networks, where $n$ is the number of neurons in a single layer of the network. See the workshop article [NO24] for more details on how this random variable is related to neural networks. $S_{n}$ is a sum of the underlying random variable $X=W\sqrt{2}\operatorname{ReLU}(G)$, which has mean zero, unit variance, $\mathbb{E}[S_{n}^{2}]=1$ and third cumulant equal to zero. The first non-zero higher cumulant is when $k=4$, which has $\kappa_{4}=\mathbb{E}[W^{4}\cdot\sqrt{2}^{4}\cdot\operatorname{ReLU}(G)^{4}]-3=3\cdot 4\cdot\frac{1}{2}\cdot 3-3=15$. Hence Theorem 6.10 gives the approximation
@@
Figure 10: Example of the 1D neural network distribution when $n=20$. We compare the pure CLT Gaussian approximation $\rho_{Guassian}$ and the approximation from (93) which labelled here as $\rho_{\text{para-Gaussian}}$, the "para-Gaussian" distribution, to mean it is a near-Gaussian approximation that includes the effect of the $1/n$ sized 4th cumulant and $H\!\!\operatorname{e}_{4}$. We also compare both to Monte Carlo samples of the random variable. The left plot shows the raw distributions, while the right plot is the difference between the theoretical approximations and the Monte Carlo estimate for the density $\rho_{\text{MonteCarlo}}$. This example is also discussed in the workshop article [NO24].
@@ 6.5 Multi-dimensional Edgeworth expansion
The Edgeworth expansion technique works in multiple dimensions for random vectors using the multi-variable Hermite polynomials $\widetilde{H}(\vec{x},V)$. This technique does not seem to be widely cited in the literature; we only find an unpublished technical note [Sob63] with the statment of this result. We believe a possible impediment to this application was dealing with the multi-variable Hermite polynomials, which are easier to work with by use of our methods in this article. In particular, the combinatorial method of Section 2 allows one to quickly calculate the Hermite polynomials even when the covariance structure is complicated.
@@
where $\text{ReLU}(x)=\max\{x,0\}$. The vector $(A_{n},B_{n})\in\mathbb{R}^{2}$ is the joint distribution of 2 outputs from a single layer of a neural network with $n$ neurons when the inputs are correlated with correlation $\cos(\theta)$. Understanding how neural networks behave for correlated inputs is important because of the depth degeneracy problem whereby the correlation tends to $1$ as the depth of the network increases, see [JN24].
@@
Explicit formulas for the $J$ functions have been calculated in [JN24]. The relevant formulas needed for the Edgeworth expansion are
@@
Figure 11: Comparing the error sizes of the pure Gaussian debsity approximation and the “para-Gaussian” density approximation, which includes Hermite polynomials of order 4 in (106) from Example 6.14. For this experiment, $\theta=0.55$, $n=20$, and $\rho_{MonteCarlo}$ is estimated from $2^{18}$ Monte Carlo samples by using Kernel Density Estimation. As a single number encapsulation of how accurate the approximation is, we also compute the KL divergence $D_{KL}=D_{KL}(\rho_{Approx},\rho_{MonteCarlo})$ from the approximation to the Monte Carlo estimate. This example is discussed in more detail in the workshop article [NO24].
@@ 7.2 Brownian martingales from the Hermite polynomials
where $X$ and $B$ are independent Brownian motions, so that $Z_{t}=B_{t}+iX_{t}$ is a complex Brownian motion. By Theorem 7.18 of [Le ̵16], holomorphic functions of conformal martingales are martingales. Therefore $Z_{t}^{n}$ is a martingale and so is $M^{(n)}_{t}$. ∎
@@
[…]
@@ 7.4 The expected characteristic function of the GUE random matrix
For the proof it will be helpful to define some function. Let
@@
Appendix B Proofs for the Dyson Brownian motion construction
B.1 Proof of Proposition 5.4
@@ Appendix B Proofs for the Dyson Brownian motion construction
The proof of Lemma B.1 relies on the Karlin–Macgregor formula [KM59]:
Theorem B.2.
@@
We now apply Lemma B.1 to each factor and obtain
@@
Remark B.3.
In Proposition 5.4 we applied Lemma B.1 with $\vec{x}=\vec{0}^{\epsilon}$, but the Lemma is only proven for fixed $\vec{x}$ (not $\vec{x}$ depending on $\epsilon$). It is a simple exercise to adapt the proof of the lemma to compute the asymptotic form of $\mathbb{P}\left(\left\{\vec{X}(t)=\vec{0}^{\epsilon}\right\}\cap NI_{t}|{\vec{X}(0)=\vec{0}^{\epsilon}}\right)$ .
B.2 Determinantal formula

### v2

@@ 7.1 Fourier formula for the Hermite polynomials
The following result corresponds to equation (6.5) in [Sobel-tech-1963].
@@ 7.4 Edgeworth expansion in one dimension
Finally, by the technical assumption that $\left|\phi\right|^{v}$ is integrable, we can use the characteristic/Fourier inversion formula for the density $\rho_{S}(x)=\frac{1}{2\pi}\intop_{-\infty}^{\infty}e^{-\imath ux}\phi_{S_{n}}(u)\mathrm{d}u$ to recover the density from this (see [feller1971probability2] Chapter XV.3 for a detailed proof of this inversion formula). By virtue of the formula for the Hermite polynomials, Proposition 7.4, we can exactly integrate the terms of the approximation for $\phi_{S}$ in (116), as follows:
@@
To complete the proof, it remains only to argue that the $O(\frac{1}{\sqrt{n}^{k-1}})$ error term in the approximation for $\phi_{S}$ in (116) (which holds uniformly for $u$ in a compact set), translates to a $O(\frac{1}{\sqrt{n}^{k-1}})$ error in the density $\rho_{S}$ after integrating with the inversion formula. This error analysis is argued by using the techincal assumption that $\intop_{-\infty}^{\infty}|\phi_{S_{n}}(u)|^{v}\mathrm{d}u<\infty$ which allows one to approximate the inversion formula integral to arbitarty accuracey using $u$ in a compact set; we refer to [feller1971probability2] Chapter XVI.2 for the detailed verification of this error analysis. ∎
@@
Instead of obtaining just the leading order correction, one can obtain an entire series expansion in powers of $n$ known as the Edgeworth expansion. This is achieved by simplifying using more terms in the approximation to the characteristic function (i.e. more cumulants) and also more terms in in the exponential approximation. For example, let us take $k=3$ so that the leading order correction is $\frac{\kappa_{3}}{3!\sqrt{n}}H\!\!\operatorname{e}_{3}(\frac{x}{\sigma})$. By expanding up to order $u^{4}$ and using the higher order correction $\exp(x)=1+x+\frac{1}{2}x^{2}+O(x^{3})$ , one will find that the next order correction is $\frac{1}{n}\left(\frac{\kappa_{4}}{4!\sigma^{4}}H\!\!\operatorname{e}_{4}(\frac{x}{\sigma})+\frac{1}{2}\big(\frac{\kappa_{3}}{3!\sigma^{3}}\big)^{2}H\!\!\operatorname{e}_{6}(\frac{x}{\sigma})\right)$ where the $H\!\!\operatorname{e}_{6}$ comes from the square term in the exponential which gives an $\frac{1}{2}(\frac{1}{3!}(\imath z)^{3})^{2}$ in the series expansion for $\phi_{S}$. We refer to [feller1971probability2] Chapter XVI for a detailed exposition.
@@
The quantity $S_{n}=\frac{1}{\sqrt{n}}\sum_{j=1}^{n}W_{i}\sqrt{2}\operatorname{ReLU}\left(G_{i}\right)$ where $G_{i},W_{i}$ are iid Gaussians and $\text{ReLU}(x)=\max\{x,0\}$ arises in the theory of neural networks, where $n$ is the number of neurons in a single layer of the network. See the workshop article [nica2024improving] for more details on how this random variable is related to neural networks. $S_{n}$ is a sum of the underlying random variable $X=W\sqrt{2}\operatorname{ReLU}(G)$, which has mean zero, unit variance, $\mathbb{E}[S_{n}^{2}]=1$ and third cumulant equal to zero. The first non-zero higher cumulant is when $k=4$, which has $\kappa_{4}=\mathbb{E}[W^{4}\cdot\sqrt{2}^{4}\cdot\operatorname{ReLU}(G)^{4}]-3=3\cdot 4\cdot\frac{1}{2}\cdot 3-3=15$. Hence Theorem 7.10 gives the approximation
@@
Figure 13: Example of the 1D neural network distribution when $n=20$. We compare the pure CLT Gaussian approximation $\rho_{Guassian}$ and the approximation from (120) which labelled here as $\rho_{\text{para-Gaussian}}$, the "para-Gaussian" distribution, to mean it is a near-Gaussian approximation that includes the effect of the $1/n$ sized 4th cumulant and $H\!\!\operatorname{e}_{4}$. We also compare both to Monte Carlo samples of the random variable. The left plot shows the raw distributions, while the right plot is the difference between the theoretical approximations and the Monte Carlo estimate for the density $\rho_{\text{MonteCarlo}}$. This example is also discussed in the workshop article [nica2024improving].
@@ 7.5 Multi-dimensional Edgeworth expansion
The Edgeworth expansion technique works in multiple dimensions for random vectors using the multi-variable Hermite polynomials $\widetilde{H}(\vec{x},V)$. This technique does not seem to be widely cited in the literature; we only find an unpublished technical note [Sobel-tech-1963] with the statment of this result. We believe a possible impediment to this application was dealing with the multi-variable Hermite polynomials, which are easier to work with by use of our methods in this article. In particular, the combinatorial method of Section 2 allows one to quickly calculate the Hermite polynomials even when the covariance structure is complicated.
@@
where $\text{ReLU}(x)=\max\{x,0\}$. The vector $(A_{n},B_{n})\in\mathbb{R}^{2}$ is the joint distribution of 2 outputs from a single layer of a neural network with $n$ neurons when the inputs are correlated with correlation $\cos(\theta)$. Understanding how neural networks behave for correlated inputs is important because of the depth degeneracy problem whereby the correlation tends to $1$ as the depth of the network increases, see [DepthDegeneracy].
@@
Explicit formulas for the $J$ functions have been calculated in [DepthDegeneracy]. The relevant formulas needed for the Edgeworth expansion are
@@
Figure 14: Comparing the error sizes of the pure Gaussian debsity approximation and the “para-Gaussian” density approximation, which includes Hermite polynomials of order 4 in (133) from Example 7.14. For this experiment, $\theta=0.55$, $n=20$, and $\rho_{MonteCarlo}$ is estimated from $2^{18}$ Monte Carlo samples by using Kernel Density Estimation. As a single number encapsulation of how accurate the approximation is, we also compute the KL divergence $D_{KL}=D_{KL}(\rho_{Approx},\rho_{MonteCarlo})$ from the approximation to the Monte Carlo estimate. This example is discussed in more detail in the workshop article [nica2024improving].
@@ 8.2 Brownian martingales from the Hermite polynomials
where $X$ and $B$ are independent Brownian motions, so that $Z_{t}=B_{t}+iX_{t}$ is a complex Brownian motion. By Theorem 7.18 of [LeGall], holomorphic functions of conformal martingales are martingales. Therefore $Z_{t}^{n}$ is a martingale and so is $M^{(n)}_{t}$. ∎
@@
This is yet another way to derive the sum formula $H\!\!\operatorname{e}_{n}(x,t)=\sum_{k=0}^{\lfloor{n/2}\rfloor}\frac{n!}{2^{k}\cdot k!\cdot(n-2k)!}(-1)^{k}t^{k}x^{n-2k}$, which is equivalent to the combinatorial formula from Section 2 .
@@ 8.4 The expected characteristic function of the GUE random matrix
For the proof it will be helpful to define some functions. Let
@@
Appendix B Proof of the Mills inequality for the Gaussian integral
We first observe that
$\intop_{A}^{\infty}e^{-B(x-c)^{2}}\mathrm{d}x$ $=\frac{1}{\sqrt{2B}}\intop_{\sqrt{2B}(A-c)}^{\infty}e^{-t^{2}/2}\,\mathrm{d}t$
which follows by making the change of variables $t=\sqrt{2B}(x-c)$, so $\mathrm{d}x=\frac{\mathrm{d}t}{\sqrt{B}}$. Therefore it suffices to prove the case when $A=\lambda$, $B=\frac{1}{2}$ and $c=0$, namely,
$\frac{1}{\lambda+\frac{1}{\lambda}}e^{-\frac{1}{2}\lambda^{2}}\leq\intop_{\lambda}^{\infty}e^{-\frac{1}{2}x^{2}}\mathrm{d}x\leq\frac{1}{\lambda}e^{-\frac{1}{2}\lambda^{2}},$
which is the form in which the Mills inequality is usually stated.
The upper bound is found by using the estimate $t/\lambda>1$ in the range of integration:
$\intop_{\lambda}^{\infty}e^{-t^{2}/2}dy\leq\intop_{\lambda}^{\infty}\frac{t}{\lambda}e^{-t^{2}/2}dy=\frac{1}{\lambda}\left[-e^{-t^{2}/2}\right]_{\lambda}^{\infty}=\frac{1}{\lambda}e^{-\lambda^{2}/2}$
To prove the lower bound, we use a remarkable anti-derivative [McKean2012]. First note that ${t^{4}+2t^{2}-1}<{t^{4}+2t^{2}+1}$ for any $y\in\mathbb{R}$. Therefore
$\intop_{\lambda}^{\infty}e^{-t^{2}/2}dy\geq\intop_{\lambda}^{\infty}e^{-t^{2}/2}\left(\frac{t^{4}+2t^{2}-1}{t^{4}+2t^{2}+1}\right)dy.$
Next comes the remarkable anti-derivative: a direct calculation shows that
$\frac{\mathrm{d}}{\mathrm{d}t}\left(-\frac{t}{t^{2}+1}e^{-t^{2}/2}\right)$ $=\frac{t^{4}+2t^{2}-1}{t^{4}+2t^{2}+1}\,e^{-t^{2}/2}.$
It follows that
$\intop_{\lambda}^{\infty}e^{-t^{2}/2}dy\geq\left[-\frac{t}{t^{2}+1}e^{-t^{2}/2}\right]_{\lambda}^{\infty}=\frac{\lambda}{\lambda^{2}+1}e^{-\lambda^{2}/2}=\frac{1}{\lambda+\frac{1}{\lambda}}e^{-\lambda^{2}/2}.$
Appendix C Extra material for the Dyson Brownian motion construction
C.1 Proof of Proposition 5.4
@@ Appendix C Extra material for the Dyson Brownian motion construction
The proof of Lemma C.1 relies on the Karlin–Macgregor formula [KarlinMcGregor]:
Theorem C.2.
@@
We now apply Lemma C.1 to each factor and obtain
@@
Remark C.3.
In Proposition 5.4 we applied Lemma C.1 with $\vec{x}=\vec{0}^{\epsilon}$, but the Lemma is only proven for fixed $\vec{x}$ (not $\vec{x}$ depending on $\epsilon$). It is a simple exercise to adapt the proof of the lemma to compute the asymptotic form of $\mathbb{P}\left(\left\{\vec{X}(t)=\vec{0}^{\epsilon}\right\}\cap NI_{t}|{\vec{X}(0)=\vec{0}^{\epsilon}}\right)$ .
C.2 Determinantal formula
@@
C.3 Proof of Proposition 6.3
Proceeding similarly to the ‘unspiked’ case, i.e. by taking the limit of $\epsilon$-of-room watermelons (see also [Johansson2001CMP]), the spiked DBM process is a biorthogonal ensemble of the form (77) with $\tilde{\psi}_{j}(x)=x^{j}$ and
$\tilde{\phi}_{j}(x)$ $=\begin{cases}x^{j}\varphi_{t}(x)\quad&\text{if }j<n\\ \varphi_{t}(x-c)&\text{if }j=n.\end{cases}$
recalling the notation $\varphi_{t}$ for the Gaussian density of variance $t$. By applying row and column operations, we can replace $\phi_{j}$ by any monic polynomial of degree $j$. We will choose $\phi_{j}=\widetilde{H}_{j}(\cdot,t)$ for some $t$ to be determined later. In order to apply Theorem 5.13, we need to compute the Gram matrix. As for the unspiked case, $G_{j\ell}=$ whenever whenever $j\neq\ell$ and $j,\ell<n$. Moreover, for any $j\in[n]$, by Proposition 8.1,
$G_{Nj}$ $=\intop_{-\infty}^{\infty}\phi_{t}(x-c)\widetilde{H}_{j}(x,t)\mathrm{d}x=\intop_{-\infty}^{\infty}\left(\sum_{k=0}^{\infty}\widetilde{H}_{k}(x,t)\left(\frac{c}{t}\right)^{k}\frac{1}{k!}\right)\widetilde{H}_{j}(x,t)\phi_{t}(x)\mathrm{d}x$
$=\sum_{k=0}^{\infty}\frac{1}{k!}\left(\frac{c}{t}\right)^{k}\intop_{-\infty}^{\infty}\widetilde{H}_{k}(x,t)\widetilde{H}_{j}(x,t)\phi_{t}(x)\mathrm{d}x$
$=\left(\frac{c}{t}\right)^{j}\frac{1}{j!}\left(t^{j}j!\right)=c^{j}.$
Using the inversion formula for block diagonal matrices, we obtain
$G^{-1}=\left[\begin{array}[]{ccccc}&&&\vdots\\ &diag\left(\frac{1}{t^{i}i!}\right)&&\vdots&\vec{0}\\ &&&\vdots\\ \ldots&\ldots&\ldots&.\\ &-c^{-n}\left(\frac{c}{t}\right)^{j}\frac{1}{j!}&&&\frac{1}{c^{n}}\end{array}\right]$
It follows that the spiked DBM is a DPP with kernel
$\widetilde{K^{(c)}}_{n,t}(x,y)$ $=\sum_{i,j=0}^{n}\tilde{\phi}_{i}(x)\tilde{\psi}_{j}(y)\left[\tilde{G}^{-t}\right]_{ij}$
$=\sum_{i=0}^{n-1}\tilde{\phi}_{i}(x)\tilde{\psi}_{i}(y)\frac{1}{t^{i}i!}+\frac{1}{c^{n}}\tilde{\phi}_{n}(x)\tilde{\psi}_{n}(y)-\frac{1}{c^{n}}\tilde{\psi}_{n}(y)\sum_{j=0}^{n-1}\left(\frac{c}{t}\right)^{j}\frac{1}{j!}\tilde{\phi}_{i}(x)$
$=\sum_{i=0}^{n-1}\frac{\widetilde{H}_{i}(x,t)\widetilde{H}_{i}(y,t)}{t^{i}i!}\varphi_{t}(x)+\frac{1}{c^{N}}\widetilde{H}_{n}(y,t)\varphi_{t}(x-c)-\frac{1}{c^{N}}p_{t}^{N}(y)\sum_{j=0}^{N-1}\left(\frac{c}{t}\right)^{j}\frac{1}{j!}p_{t}^{j}(x)\varphi_{t}(x).$
Proposition 6.3 now follows by conjugating $\widetilde{K^{(c)}}$ with the function $\sqrt{\phi_{t}}$.
C.4 Proof of (96)
We have
$\sum_{j=0}^{n-1}c^{j-n}\left(\frac{n}{\sigma}\right)^{(j-n)/2}\sqrt{\frac{n}{\sigma}}$ $\int_{\gamma-\epsilon}^{\infty}\left|\Psi_{n}\left(\frac{x\sqrt{n}}{\sqrt{\sigma}}\right)\right|\left|\Psi_{j}\left(\frac{x\sqrt{n}}{\sqrt{\sigma}}\right)\right|\,dx$
$\leq\sum_{j=0}^{n-1}c^{j-n}\left(\frac{n}{\sigma}\right)^{(j-n)/2}\sqrt{\frac{n}{\sigma}}\int_{\gamma-\epsilon}^{\infty}\frac{C_{\delta}}{(nj)^{1/4}}e^{-(n+j)(x/(2\sqrt{\sigma})-1)^{2}/2}\,dx$
$\leq\sqrt{\frac{n}{\sigma}}\sum_{j=0}^{n-1}\left(\frac{c^{2}n}{\sigma}\right)^{(j-n)/2}\frac{C_{\delta}}{(nj)^{1/4}}\frac{1}{(n+j)(\gamma-\epsilon-2\sqrt{\sigma})}e^{-(n+j)\frac{\gamma-\epsilon}{2\sqrt{\sigma}}/2}$
$\leq n^{-3/4}\frac{C_{\delta}}{\sqrt{\sigma}\delta}\sum_{j=0}^{n-1}\left(\frac{c^{2}n}{\sigma}\right)^{(j-n)/2}e^{-(n+j)(1+\frac{\delta}{2\sqrt{\sigma}})/2}$
$=n^{-3/4}\frac{C_{\delta}}{\sqrt{\sigma}\delta}e^{-n(1+\frac{\delta}{2\sqrt{\sigma}})/2}\left(\frac{c^{2}n}{\sigma}\right)^{-n/2}\sum_{j=0}^{n-1}\left(\sqrt{\frac{c^{2}n}{\sigma}}e^{-(1+\frac{\delta}{2\sqrt{\sigma}})/2}\right)^{j}$
$\leq Ce^{-n(1+\frac{\delta}{2\sqrt{\sigma}})/2}$
as claimed.
C.5 Proof of (99)
For $x>\gamma-\epsilon>2\sqrt{\sigma}$,
$f^{\prime}(x)$ $=\frac{1}{2\sqrt{\sigma}}S^{\prime}\left(\frac{x}{2\sqrt{\sigma}}\right)-\frac{x-c}{\sigma}$
$=\frac{1}{\sqrt{\sigma}}\left(\frac{x}{2\sqrt{\sigma}}-\sqrt{\left(\frac{x}{2\sqrt{\sigma}}\right)^{2}-1}\right)-\frac{x-c}{\sigma}$
$=\frac{2c-x}{2\sigma}-\frac{1}{2\sigma}\sqrt{x^{2}-4\sigma}.$
Thus $f^{\prime}(x)=0$ if and only if $2c-x=\sqrt{x^{2}-4\sigma}$ iff $4c^{2}-4cx+x^{2}=x^{2}-4\sigma$. Thus, the only critical point is $x_{\ast}=c+\frac{\sigma}{c}$.